Product Profitability Analysis

PowerTrain Sports Inc. manufactures and sells two styles of All Terrain Vehicles (ATVs), the Mountain Monster and Desert Dragon, from a single manufacturing facility. The manufacturing facility operates at 100% of capacity. The following per-unit information is available for the two products:

Mountain Monster Desert Dragon
Sales price $6,200 $4,000
Variable cost of goods sold 3,910 2,680
Manufacturing margin $2,290 $1,320
Variable selling expenses 1,360 480
Contribution margin $930 $840
Fixed expenses 440 340
Income from operations $490 $500

In addition, the following sales unit volume information for the period is as follows:

Mountain Monster Desert Dragon
Sales unit volume 2,800 2,000

a. Prepare a contribution margin by product report. Calculate the contribution margin ratio for each product as a whole percent.

PowerTrain Sports Inc.
Contribution Margin by Product
Mountain Monster Desert Dragon
$ $
$ $
$ $
% %
Answer #1
Mountain Monster Desert Dragon
Revenues 17360000 8000000
Variable Cost of goods sold 10948000 5360000
Manufacturing margin 6412000 2640000
Variable selling expenses 3808000 960000
Contribution margin 2604000 1680000
Contribution margin ratio 15% 21%
Formulas used:
Revenues =2800*6200 =2000*4000
Variable Cost of goods sold =2800*3910 =2000*2680
Variable selling expenses =2800*1360 =2000*480